OpenStudy (anonymous):

tan(60)=SQUARE ROOT 3=21/x

OpenStudy (campbell_st):

so that means if \[\tan(60) = \frac{21}{x}\] then \[x = \frac{21}{\tan(60)}\] using the substitution for the exact value \[x = \frac{21}{\sqrt{3}}\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Can i simplify that

OpenStudy (campbell_st):

you may need to rational the denominator \[\frac{21}{\sqrt{3}} \times \frac{\sqrt{3}}{\sqrt{3}} = \frac{21 \sqrt{3}}{3}\] which can be simplified.